OpenStudy (anonymous):

Put these in y=mx+b form: 3x+2y-1=0 x-2y+13=0

OpenStudy (anonymous):

First, what system of equation is this, substitution or elimination?

OpenStudy (anonymous):

It just says to graph the equations, so I think I have to put it in y=mx+b form to graph? or is that not right

OpenStudy (shinalcantara):

since you're asked to have it in slope intercept form, then transpose the terms x and constant to the right side of the equation leaving the term with y on at the left. divide the whole equation with the coefficient of y so that y will have 1 as it coefficient. now try to solve it

OpenStudy (anonymous):

I got y=-3/2x+1 and y=1/2-6.5 Is that correct?

OpenStudy (shinalcantara):

try to recheck the constant for the 1st equation and the sign for the 2nd